OpenStudy (anonymous):

is the line through points P(0,5) and Q(-1,8) parallel to the line through points R(3,3) and S(5,-1)?

OpenStudy (anonymous):

y-y1 = m(x-x1) so just pick a point to use (-1,1) find slope m = (1+3)/(-1+10) = 4/9 y -1=4/9(x+1) y = 4/9x + 4/9 + 1 y = 4/9 x + 13/9